(Can someone please help me with this question?)
The residents of a downtown neighborhood designed a triangular-shaped park as part of a city beautification program. The park is bound by streets on all sides. The second angle of the triangle is 5^o more than the first. The third angle is 5^o less than eight times the first. Find the measures of the angles.
The measure of the first angle is x=?
The measure of the second angle is x+5=?
The measure of the third angle is 8x-5=?

Answer by richwmiller(17219)

The key to the problem is that all the angles of a triangle add up to 180.
x+x+5+8x-5=180
x=18
x+5=23
18+23=41
180-41=139
or
8*18-5=139
18, 23, 139